# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v1

@@ 6.1 Toy example
Method Grid Random TPE MPCC MEHA AGILS
Time(s) 2.14 25.72 98.92 990.91 0.42 0.14
Error 0.71 0.79 0.76 0.27 0.00 0.00
@@
Finally, we evaluate the scalability of AGILS by testing it on toy examples with varying problem dimensions. The parameters of AGILS are as specified earlier, except that $p_{0}$ is set as 1. AGILS terminates when $\text{Error}<1/n$. The results in Figure 2 show that computational time increases steadily as problem dimension grows. This demonstrates that AGILS remains both efficient and stable as the problem dimension increases, highlighting its ability to handle large-scale problems effectively.
@@ 6.2 Sparse group Lasso
${\rm s.t.}$ $y\in\underset{\hat{y}\in{{\mathbb{R}}}^{m}}{\mathrm{argmin}}\,\frac{1}{2|I_{\text{tr}}|}\sum_{i\in I_{\text{tr}}}\lvert b_{i}-\hat{y}^{T}a_{i}\rvert^{2}+\sum_{j=1}^{J}x_{j}\lVert\hat{y}^{\left(m\right)}\rVert_{2}+x_{J+1}\lVert\hat{y}\rVert_{1},$
@@
where $I_{\text{val}}$ and $I_{\text{tr}}$ are the sets of validation and training sample indices, respectively. In this part, we compare the performance of our AGILS algorithm with several hyperparameter optimization methods, including grid search, random search, TPE, IGJO, VF-iDCA, and MEHA. The MPCC approach is excluded from the comparison due to its prohibitively high computational runtime. Following [32], the data generation process is as follows: $a_{i}\in\mathbb{R}^{m}$ is sampled from $N(0,I)$, and the response $b_{i}$ is computed as $b_{i}=y^{T}a_{i}+\sigma\epsilon_{i}$, where noise $\epsilon$ is sampled from $N(0,1)$. The coefficient vector $y$ is divided equally into five groups, with the $i$th group having the first $2i$ features set to value $2i$ and the remaining features set to zero, for $i=1,\ldots,5$. The noise level $\sigma$ is chosen to ensure $\text{SNR}=\|Ay\|/\|b-Ay\|=3$. The dataset is then randomly split into training, validation, and test sets of sizes $n_{\text{tr}}$, $n_{\text{val}}$, and $n_{\text{test}}$, respectively.
@@
For grid search, random search and TPE, we parameterize $x_{i}$ with the log-transformed variables $\rho_{i}=\log_{10}(x_{i})$ and the search range is $[-9,2]^{J+1}$. For grid search, we set $\rho_{1}=\rho_{2}=\dots=\rho_{M}$ and use a $20\times 20$ grid. Random search samples 400 points uniformly within the range. For TPE, the search space is modeled as a uniform distribution over the same range. For VF-iDCA, MEHA, IGJO, and AGILS, the initial value of $x_{0}$ is set to $[1,\dots,1]$. Both $y_{0}$ and $\theta_{0}$ for MEHA and AGILS are initialized as $[1,\dots,1]$. For VF-iDCA, the parameters are selected as $\beta_{0}=5$, $\rho=0.01$, $c=1$, and $\delta=0.5$. VF-iDCA terminates when $\max\{\lVert{z}^{k+1}-{z}^{k}\rVert/\sqrt{1+\lVert{z}^{k}\rVert^{2}},\tilde{t}^{k+1}/m\}<0.1$, where $\tilde{t}^{k+1}$ is defined in [28]. The IGJO algorithm is run with a maximum of 50 iterations and a minimum step size of $10^{-6}$. For MEHA, the parameters are searched over $\tilde{c}_{0}\in\{1,10,20\}$ and $p\in\{0.1,0.2,0.3,0.4\}$, while $\tilde{\alpha}$, $\tilde{\beta}$, $\tilde{\eta}$, and $\tilde{\gamma}$ scaled proportionally from the AGILS parameters ${\alpha}$, ${\beta}$, ${\eta}$, and ${\gamma}$ using scales $\{1,1/2,1/4,1/8\}$. The optimal parameters that minimize the validation error are selected as $\tilde{c}_{0}=20$, $p=0.1$, $\tilde{\gamma}=\gamma$, $\tilde{\alpha}=\alpha$, $\tilde{\beta}=\beta/9$, and $\tilde{\eta}=\eta/5$. MEHA terminates when $\lVert\mathbf{z}^{k+1}-\mathbf{z}^{k}\rVert/\sqrt{1+\lVert\mathbf{z}^{k}\rVert^{2}}<0.005/m$. AGILS uses the same parameter settings as in the toy example, with specific values for constants such as $L_{F_{x}}=L_{f_{x}}=\rho_{f_{1}}=\rho_{f_{2}}=L_{g_{1}}=0$, $L_{F_{y}}=\lambda_{\text{max}}(A_{\text{val}}^{\top}A_{\text{val}})/n_{\text{val}}$, $L_{f_{y}}=\lambda_{\text{max}}(A_{\text{tr}}^{\top}A_{\text{tr}})/n_{\text{tr}}$, $\rho_{g_{1}}=1$, $\rho_{g_{2}}=m$. The penalty parameters are set as $p_{0}=6$, $\varrho_{p}=0.01$. AGILS terminates when $\lVert z^{k+1}-z^{k}\rVert/\sqrt{1+\lVert z^{k}\rVert^{2}}<0.005/m$, $z^{k}:=(x^{k},y^{k})$ and $t^{k+1}<0.1$. Proximal gradient method is used to compute $\theta^{k}$ that satisfies inexact criteria (18) and (21).
@@
The experiments are repeated 20 times with $n_{\text{tr}}=200$, $n_{\text{val}}=200$, $n_{\text{test}}=200$, and $m=300$. Results are reported in Table 2 and Figure 3. We report the validation error (“Val. Err.”) and test error (“Test. Err.”), computed as $\sum_{i\in I_{\text{val}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}/|I_{\text{val}}|$ and $\sum_{i\in I_{\text{test}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}/|I_{\text{test}}|$, respectively. Both errors are evaluated using the sparse group Lasso estimator, which is obtained by solving the lower-level problem with the hyperparameter value corresponding to the iterates produced by the algorithms. Additionally, we report a second type of test error, referred to as the infeasible test error (“Test. Err. Infeas.”), where the lower-level variable values are taken directly from the iterates generated by VF-iDCA, MEHA and AGILS. “Feasibility” is a scaled measure of the value function constraint violation, given by $(\varphi(x^{k},y^{k})-v(x^{k}))/|I_{\text{val}}|$ for VF-iDCA, and $(\varphi(x^{k},y^{k})-v_{\gamma}(x^{k},y^{k}))/|I_{\text{val}}|$ for MEHA and AGILS.
@@
Table 2: Comparison of different methods on the sparse group Lasso bilevel hyperparameter selection problem with $n_{\text{tr}}=n_{\text{val}}=n_{\text{test}}=200,m=300$
@@
Grid 50.91(1.79) 168.26(23.91) 170.10(21.28) - -
Random 81.53(3.76) 185.87(27.24) 193.93(26.39) - -
TPE 137.73(4.45) 168.16(21.33) 178.96(22.70) - -
IGJO 192.95(99.66) 139.43(12.32) 169.62(21.11) - -
VF-iDCA 88.59(37.66) 134.68(14.71) 150.40(12.02) 138.70(11.77) 0.04(0.03)
MEHA 16.18(3.77) 98.35(8.79) 158.61(13.62) 158.42(13.80) 0.00(0.00)
AGILS 14.93(4.31) 96.06(10.53) 155.25(12.48) 155.39(12.05) 0.00(0.00)
@@
As shown in Table 2, our proposed method, AGILS, outperforms the others in terms of validation error while requiring the least computation time. The validation error, corresponding to the value of the upper-level objective at a feasible point derived from the hyperparameter value of the iterates, can be seen as an indicator of optimization performance. AGILS also achieves competitive test errors. Interestingly, VF-iDCA finds the parameter $y$ that has the lowest test error, but the corresponding constraint violation is large. MEHA performs similarly to AGILS, but slightly less effective. Moreover, MEHA’s performance is highly sensitive to algorithmic parameters, and selecting inappropriate parameters can significantly degrade its effectiveness. Figure 3 illustrates that AGILS has fast convergence during the initial iterations.
Next, we investigate the impact of different solvers for solving the proximal lower-level problem in AGILS. We compare three algorithms: the Proximal Gradient Method (PGM), the Fast Iterative Shrinkage-Thresholding Algorithm (FISTA), and the Alternating Direction Method of Multipliers (ADMM). Each experiment is repeated 20 times, with all other parameters of AGILS held consistent with those used above. The results, presented in Table 3, show that the AGILS framework is robust to different solvers, with each solver exhibiting similar performance. This supports the flexibility and adaptability of the AGILS framework.
Table 3: Comparison of AGILS with different proximal lower-level problem solvers on the sparse group Lasso bilevel hyperparameter selection problem with $n_{\text{tr}}=n_{\text{val}}=n_{\text{test}}=200,m=300$
@@
AGILS(PGM) 14.93(4.31) 96.06(10.53) 155.25(12.48) 155.39(12.05) 0.00(0.00)
AGILS(FISTA) 15.82(5.07) 96.06(10.53) 155.25(12.48) 155.39(12.05) 0.00(0.00)
AGILS(ADMM) 15.15(4.38) 95.60(10.11) 154.27(12.30) 154.58(12.11) 0.00(0.00)
@@
Table 4: AGILS on large-scale sparse group Lasso bilevel hyperparameter selection problems
@@
$n_{\text{tr}}=1000,\ n_{\text{val}}=1000$  $\ n_{\text{test}}=1000,\ m=1500$ 20.13(3.72) 113.07(5.87) 116.85(9.54) 114.33(9.08) 0.00(0.00)
$n_{\text{tr}}=3000,\ n_{\text{val}}=3000$  $\ n_{\text{test}}=3000,\ m=4500$ 131.14(16.12) 106.91(2.92) 112.13(3.06) 109.95(2.70) 0.00(0.00)
$n_{\text{tr}}=5000,\ n_{\text{val}}=5000$  $\ n_{\text{test}}=5000,\ m=7500$ 433.96(35.61) 107.95(2.45) 106.64(1.50) 105.28(1.51) 0.00(0.00)
$n_{\text{tr}}=7000,\ n_{\text{val}}=7000$  $\ n_{\text{test}}=7000,\ m=10500$ 839.55(106.05) 105.71(1.80) 105.91(2.15) 104.72(2.14) 0.00(0.00)

### v2

@@ 6.1 Toy example
Method Grid Random TPE MPCC IGJO VF-iDCA MEHA AGILS
Orig Abs Rel
Time(s) 2.55 27.97 107.37 179.12 55.19 2.27 0.48 0.14 0.15 0.39
Error 0.71 0.79 0.76 1.00 0.69 0.10 0.00 0.00 0.00 0.00
We report the average inner-to-outer iteration ratio, defined as $\sum_{k=1}^{K}\ell_{k}/K$, where $\ell_{k}$ is the number of inner iterations at the $k$-th outer step. Table 2 shows the overall average and the averages over the first 200 and remaining outer iterations. Inner iterations tend to be more variable in the early phase and more stable later.
Table 2: Average inner-to-outer iteration ratio of AGILS and its variants for different problem dimensions
Method Dimension $n=200$ Dimension $n=600$
Overall First 200 After 200 Overall First 200 After 200
AGILS 2.84 3.31 2.46 2.87 5.29 2.03
AGILS_A 3.38 4.45 2.46 2.88 5.29 2.03
AGILS_R 6.28 7.48 4.20 43.39 64.59 3.00
@@
Finally, we evaluate the scalability of AGILS by testing it on toy examples with varying problem dimensions. The parameters of AGILS are as specified earlier, except that $p_{0}$ is set as 1. AGILS terminates when $\text{Error}<1/n$. The results in Figure 2 show that computational time increases steadily as problem dimension grows. This demonstrates that AGILS remains both efficient and stable as the problem dimension increases, highlighting its ability to handle large-scale problems effectively. Furthermore, in all repeated runs of the toy example, the feasibility correction procedure, designed to enforce constraint satisfaction, was never triggered. This observation is also reflected in the reported Error.
@@ 6.2 Sparse group Lasso
${\rm s.t.}$ $y\in\underset{\hat{y}\in{{\mathbb{R}}}^{m}}{\mathrm{argmin}}\,\frac{1}{2|I_{\text{tr}}|}\sum_{i\in I_{\text{tr}}}\lvert b_{i}-\hat{y}^{T}a_{i}\rvert^{2}+\sum_{j=1}^{J}x_{j}\lVert\hat{y}^{\left(j\right)}\rVert_{2}+x_{J+1}\lVert\hat{y}\rVert_{1},$
@@
where $I_{\text{val}}$ and $I_{\text{tr}}$ are the sets of validation and training sample indices, respectively. In this part, we compare the performance of our AGILS and its variants with several hyperparameter optimization methods, including grid search, random search, TPE, IGJO, VF-iDCA, and MEHA. The MPCC approach is excluded from the comparison due to its prohibitively high computational runtime. Following [26], the data generation process is as follows: $a_{i}\in\mathbb{R}^{m}$ is sampled from $N(0,I)$, and the response $b_{i}$ is computed as $b_{i}=y^{T}a_{i}+\sigma\epsilon_{i}$, where noise $\epsilon$ is sampled from $N(0,I)$. The coefficient vector $y$ is divided equally into five groups, with the $i$th group having the first $2i$ features set to value $2i$ and the remaining features set to zero, for $i=1,\ldots,5$. The noise level $\sigma$ is chosen to ensure $\text{SNR}=\|Ay\|/\|b-Ay\|=3$. The dataset is then randomly split into training, validation, and test sets of sizes $n_{\text{tr}}$, $n_{\text{val}}$, and $n_{\text{test}}$, respectively.
@@
For grid search, random search and TPE, we parameterize $x_{i}$ with the log-transformed variables $\rho_{i}=\log_{10}(x_{i})$ and the search range is $[-9,2]^{J+1}$. For grid search, we set $\rho_{1}=\rho_{2}=\dots=\rho_{M}$ and use a $20\times 20$ grid. Random search samples 400 points uniformly within the range. For TPE, the search space is modeled as a uniform distribution over the same range. For VF-iDCA, MEHA, IGJO, AGILS and its variants, the initial value of $x_{0}$ is set to $[1,\dots,1]$. Both $y_{0}$ and $\theta_{0}$ for MEHA, AGILS and its two variants are initialized as $[1,\dots,1]$. For VF-iDCA, the parameters are selected as $\beta_{0}=5$, $\rho=0.01$, $c=1$, and $\delta=0.1$. VF-iDCA terminates when $\max\{\lVert{z}^{k+1}-{z}^{k}\rVert/\sqrt{1+\lVert{z}^{k}\rVert^{2}},\tilde{t}^{k+1}/m\}<0.1$, where $\tilde{t}^{k+1}$ is defined in [30]. The IGJO algorithm is run with a maximum of 50 iterations and a minimum step size of $10^{-6}$. For MEHA, the parameters are searched over $\tilde{c}_{0}\in\{1,10,20\}$ and $p\in\{0.1,0.2,0.3,0.4\}$, while $\tilde{\alpha}$, $\tilde{\beta}$, $\tilde{\eta}$, and $\tilde{\gamma}$ scaled proportionally from the AGILS parameters ${\alpha}$, ${\beta}$, ${\eta}$, and ${\gamma}$ using scales $\{1,1/2,1/4,1/8\}$. The optimal parameters that minimize the validation error are selected as $\tilde{c}_{0}=20$, $p=0.1$, $\tilde{\gamma}=\gamma$, $\tilde{\alpha}=\alpha$, $\tilde{\beta}=\beta/8$, and $\tilde{\eta}=\eta/4$. MEHA terminates when $\lVert\mathbf{z}^{k+1}-\mathbf{z}^{k}\rVert/\sqrt{1+\lVert\mathbf{z}^{k}\rVert^{2}}<0.005/m$. AGILS and its variants use the same parameter settings as in the toy example, except for the penalty parameters and the following constants such as $L_{F_{x}}=L_{f_{x}}=\rho_{f_{1}}=\rho_{f_{2}}=L_{g_{1}}=0$, $L_{F_{y}}=\lambda_{\text{max}}(A_{\text{val}}^{\top}A_{\text{val}})/n_{\text{val}}$, $L_{f_{y}}=\lambda_{\text{max}}(A_{\text{tr}}^{\top}A_{\text{tr}})/n_{\text{tr}}$, $\rho_{g_{1}}=1$, $\rho_{g_{2}}=m$. The penalty parameters are set as $p_{0}=6$, $\varrho_{p}=0.01$. AGILS and its two variants terminate when $\lVert z^{k+1}-z^{k}\rVert/\sqrt{1+\lVert z^{k}\rVert^{2}}<0.005/m$, $z^{k}:=(x^{k},y^{k})$ and $t^{k+1}<0.1$. Proximal gradient method is used to compute $\theta^{k}$ that satisfies inexact criteria (17) and (20). Following the same algorithmic structure, AGILS and its variants adopt a decaying absolute tolerance $s_{k}=5/(k+1)^{1.05}$ and a relative criterion with parameter $\tau_{k}=10/(k+1)^{0.2}$.
@@
The experiments are repeated 20 times with $n_{\text{tr}}=200$, $n_{\text{val}}=200$, $n_{\text{test}}=200$, and $m=300$. Results are reported in Table 3 and Figure 3 correspond to the mean values computed over 20 independent runs, with standard deviations shown in parentheses. We report the validation error (“Val. Err.”) and test error (“Test. Err.”), computed as $\sum_{i\in I_{\text{val}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}/|I_{\text{val}}|$ and $\sum_{i\in I_{\text{test}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}/|I_{\text{test}}|$, respectively. Both errors are evaluated using the sparse group Lasso estimator, which is obtained by solving the lower-level problem with the hyperparameter value $x^{k}$ produced by the algorithms. Additionally, we report a second type of test error, referred to as the infeasible test error (“Test. Err. Infeas.”) $\sum_{i\in I_{\text{test}}}\left|b_{i}-(y^{k})^{\top}a_{i}\right|^{2}/|I_{\text{test}}|$, where the lower-level variable values are taken directly from the iterates generated by VF-iDCA, MEHA, AGILS and its variants. “Feasibility” is a scaled measure of the value function constraint violation, given by $(\varphi(x^{k},y^{k})-v(x^{k}))/|I_{\text{val}}|$ for VF-iDCA, and $(\varphi(x^{k},y^{k})-v_{\gamma}(x^{k},y^{k}))/|I_{\text{val}}|$ for MEHA, AGILS and its two variants. Note that this post-processing step is not part of the algorithm and is excluded from runtime.
@@
Table 3: Comparison of different methods on the sparse group Lasso bilevel hyperparameter selection problem with $n_{\text{tr}}=n_{\text{val}}=n_{\text{test}}=200,m=300$.
@@
Grid 50.73(3.65) 168.26(23.91) 170.10(21.28) - -
Random 79.71(4.94) 185.87(27.24) 193.93(26.39) - -
TPE 135.66(7.94) 171.42(21.93) 182.19(25.32) - -
IGJO 173.73(90.01) 139.43(12.32) 169.62(21.11) - -
VF-iDCA 84.71(47.81) 136.16(32.39) 148.85(31.06) 129.46(15.47) 0.08(0.04)
MEHA 13.15(1.87) 98.41(8.83) 158.63(13.52) 158.39(13.76) 0.00(0.00)
AGILS 12.25(2.11) 95.93(9.71) 153.75(12.72) 153.14(12.24) 0.00(0.00)
AGILS_A 12.82(2.34) 96.06(10.53) 155.25(12.48) 155.39(12.05) 0.00(0.00)
AGILS_R 12.37(2.22) 95.93(9.71) 153.75(12.72) 153.13(12.24) 0.00(0.00)
@@
As shown in Table 3, our proposed method, AGILS, outperforms the others in terms of validation error while requiring the least computation time. The validation error, corresponding to the value of the upper-level objective at a feasible point derived from the hyperparameter value of the iterates, can be seen as an indicator of optimization performance. AGILS also achieves competitive test errors. Notably, both variants, AGILS_A and AGILS_R, exhibit similar performance to AGILS in terms of both validation and test errors. Interestingly, although VF-iDCA finds a parameter $y$ that achieves the lowest test error in the infeasible setting, the corresponding constraint violation is large. This suggests that VF-iDCA may possess good generalization performance, even when feasibility is not strictly enforced. MEHA performs similarly to AGILS, but slightly less effective. Moreover, MEHA’s performance heavily relies on proper selection of algorithmic parameters, and selecting inappropriate parameters can significantly degrade its effectiveness. Figure 3 illustrates that AGILS has fast convergence during the initial iterations. The average inner-to-outer iteration ratios for AGILS and its variants are approximately 2. This indicates that the two inexactness strategies yield similar computational efficiency.
Next, we investigate the impact of different solvers for solving the proximal lower-level problem in AGILS. We compare three algorithms: the Proximal Gradient Method (PGM), the Fast Iterative Shrinkage-Thresholding Algorithm (FISTA), and the Alternating Direction Method of Multipliers (ADMM). Each experiment is repeated 20 times, with all other parameters of AGILS held consistent with those used above. The results, presented in Table 4 as the mean and the standard deviation over the 20 runs, show that the AGILS framework is robust to different solvers, with each solver exhibiting similar performance. This supports the flexibility and adaptability of the AGILS framework.
Table 4: Comparison of AGILS with different proximal lower-level problem solvers on the sparse group Lasso bilevel hyperparameter selection problem with $n_{\text{tr}}=n_{\text{val}}=n_{\text{test}}=200,m=300$.
@@
AGILS(PGM) 12.25(2.11) 95.93(9.71) 153.75(12.72) 153.14(12.24) 0.00(0.00)
AGILS(FISTA) 11.85(2.04) 95.93(9.71) 153.75(12.72) 153.14(12.24) 0.00(0.00)
AGILS(ADMM) 12.13(2.06) 95.60(10.11) 154.27(12.30) 154.58(12.11) 0.00(0.00)
@@
Similarly, during all repeated runs of the sparse group Lasso experiments, the feasibility correction procedure, introduced to enforce feasibility, was never triggered indicating that infeasibility was rarely encountered.
Table 5: AGILS on large-scale sparse group Lasso bilevel hyperparameter selection problems.
@@
$n_{\text{tr}}=1000,\ n_{\text{val}}=1000$  $\ n_{\text{test}}=1000,\ m=1500$ 24.48(3.21) 112.94(5.85) 116.72(9.49) 114.23(9.05) 0.00(0.00)
$n_{\text{tr}}=3000,\ n_{\text{val}}=3000$  $\ n_{\text{test}}=3000,\ m=4500$ 133.71(11.18), 106.89(2.91) 112.11(3.06) 109.93(2.69) 0.00(0.00)
$n_{\text{tr}}=5000,\ n_{\text{val}}=5000$  $\ n_{\text{test}}=5000,\ m=7500$ 445.44(42.99) 107.94(2.45) 106.64(1.50) 105.28(1.51) 0.00(0.00)
$n_{\text{tr}}=7000,\ n_{\text{val}}=7000$  $\ n_{\text{test}}=7000,\ m=10500$ 787.01(74.59) 105.71(1.80) 105.91(2.15) 104.72(2.14) 0.00(0.00)